OpenStudy (anonymous):

graph the equation and identify the y-intercept y+x=-3

OpenStudy (anonymous):

The simplified equation is y = -x - 3 You plug in x = 0 to find the y-intercept Y-intercept = (0,-3)

OpenStudy (anonymous):

will i need to graph 0.-3